OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 7 tan3x - 21 tan x = 0 (I have no idea what to do)

OpenStudy (aum):

tan(3x) can be expanded. Are you familiar with that trig identity?

OpenStudy (aum):

\[ \tan(3x) = \frac{3\tan(x)-\tan^3(x)}{1-3\tan^2(x)} \]

OpenStudy (aum):

\[ 7\tan(3x) - 21\tan(x) = 0 \\ \text{Divide by 7:}\\ \tan(3x) - 3\tan(x) = 0 \\ \tan(3x) = 3\tan(x) \\ \frac{3\tan(x)-\tan^3(x)}{1-3\tan^2(x)} = 3\tan(x) \\ \text{Cross multiply, simplify and solve for x:} \\ 3\tan(x)-\tan^3(x) = 3\tan(x)(1-3\tan^2(x)) \\ 3\tan(x)-\tan^3(x) = 3\tan(x) - 9\tan^3(x)) \\ \text{Subtract 3 tan(x):} \\ -\tan^3(x) = - 9\tan^3(x)) \\ \text{Add 9 tan^3(x):}\\ 8\tan^3(x) = 0 \\ \tan^3(x) = 0 \\ \tan(x) = 0 \\ x = 0, ~~\pi, ~~2\pi \]